Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufactu5. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.)

6. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

7. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

8. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)
Answer #1
Molding Fabrication Total
Estimated total machine hours used 2500 1500 4000
(*) Estimated variable manufacturing overhead per machine hour 3.30 4.10
Estimated total variable manufacturing overhead 8250 6150 14400
Estimated total fixed manufacturing overhead 14750 17850 32600
Estimated total manufacturing overhead 23000 24000 47000
Plantwide predetermined overhead rate = Estimated total manufacturing overhead / Estimated total machine hours used = 47000 / 4000 = 11.75 per MH
Job P Job Q
Total actual machine hours used 6100 5500
Plantwide predetermined overhead rate 11.75 11.75
Manufacturing overhead applied 71675 64625
Job P
Direct materials 32000
Direct labor cost 36200
Manufacturing overhead applied 71675
Total manufacturing cost 139875
Required 5 :
Job Q
Direct materials 17500
Direct labor cost 15100
Manufacturing overhead applied 64625
Total manufacturing cost 97225
Required 6 :
Job Q
Total manufacturing cost 97225
(/) No. of units 30
Unit product cost 3241
Required 7 :
Job P Job Q
Total manufacturing cost 139875 97225
(+) Markup @ 80% 111900 77780
Total price for the job 251775 175005
(/) No. of units 20 30
Selling price per unit 12589 5834
Required 8 :
Cost of goods sold = Total manufacturing costs of Jobs sold = 139875 + 97225 = 237100